Of sin A=5/13 the evaluate 5 sin A- 2 cos A divided by tan A

Grade:10

1 Answers

Soumendu Majumdar
5 years ago
Dear Student,
The question doesn't clearly state what is being divided by tan A, still, I’ll try to solve it by dividing only  2cos A by tan A & the whole expression by tan A
Now
if only 2 cos A is being divided by tan A then it is 2 cos^2A/sinA
now by Pythagorean triplets, we know  5^2 + 12^2 = 13^2
so cos A = 12/13
so  cos^2 A = 144/169
Hence 5 sin A - 2cos^2A/sinA = 25/13 - 288/65 = - 163/65
Now if the whole expression is divided by tan A, then
the expression becomes 
5 cos A - 2cos^2A/sinA
60/13 - 288/65 = 12/65
Hope it helps!